# Diffusion-SDPO: Safeguarded Direct Preference Optimization for Diffusion Models

### v1

@@ 1 Introduction
Text-to-image diffusion models diffusion_survey have achieved remarkable success in generating diverse and high-quality images flux; google2025_nano_banana. However, aligning these powerful generative models with nuanced human preferences remains a critical challenge. Recent approaches have begun to incorporate human feedback rlhf into diffusion model training, drawing inspiration from alignment techniques used in large language models. In particular, Direct Preference Optimization (DPO) llm_dpo has emerged as a promising alternative to reinforcement learning for finetuning on human preferences. DPO directly optimizes the model on pairwise human comparisons (winner vs. loser outputs), and has been successfully adapted to text-to-image diffusion models in methods diffusion-dpo; mapo; dspo; dmpo; diffusion-kto to improve visual appeal and prompt alignment. Despite these advances, we find that existing DPO-based alignment of diffusion models still faces a fundamental limitation: simply maximizing the preference margin between “winner” and “loser” outputs does not necessarily translate to better absolute generation quality of the finetuned model.
@@
In our empirical analysis, we find that standard Diffusion-DPO diffusion-dpo exhibits unstable training dynamics, and the model’s generative quality can deteriorate as training proceeds. As illustrated in the left part of Fig. 1, we find that both the winner’s and loser’s denoising losses tend to increase over time, even though the preference margin ($\mathcal{L}^{w}-\mathcal{L}^{l}$) becomes more negative in the intended direction. This indicates that the model is widening the relative preference gap by making the less-preferred outputs worse, rather than truly improving the preferred outputs. In other words, relative alignment comes at the expense of absolute quality. The lack of a safeguard on the winner’s loss in existing DPO objectives leads to unstable training and potential collapse, corroborating observations in prior work dpop; calibrated-dpo; see-dpo that overly aggressive preference optimization can harm generative performance. These findings motivate the need for a new approach to preference-based diffusion finetuning that can increase preference alignment while preserving or improving the quality of the preferred outputs.
@@
To address this challenge, we propose Diffusion-SDPO [^1] Throughout the text, “Diffusion-SDPO” is used as a conceptual umbrella for our method and its guiding principles. When referring to concrete instantiations, we write “X+SDPO” to denote the integration of SDPO with a specific base DPO variant $X$ (e.g., Diffusion-DPO, DSPO, DMPO), which clarifies the application setting and configuration. – a Safeguarded Direct Preference Optimization method for diffusion models. The key idea in Diffusion-SDPO is to introduce a simple yet effective winner-preserving update rule that controls the influence of the loser sample’s gradient at each training step. In contrast to standard DPO llm_dpo; diffusion-dpo which updates the model by contrasting winner and loser equally, we derive an adaptive scaling factor for the loser’s gradient based on the geometry of the winner and loser gradients. Intuitively, our method downweights the loser branch’s contribution whenever its gradient is misaligned with the winner’s gradient. Grounded in a first-order analysis, the safeguard computes a closed-form $\lambda_{\text{safe}}$ from the inner product of the winner and loser gradients, guaranteeing that each step does not worsen the preferred output’s reconstruction loss. In practice, Diffusion-SDPO seamlessly modifies the DPO objective with this adaptive loser scaling (see Fig. 1, right), which expands the preference margin while strictly controlling the absolute error of preferred outputs. Notably, our approach is model-agnostic and can be applied on top of various diffusion alignment frameworks diffusion-dpo; mapo; dmpo; dspo, acting as a plug-in optimizer that stabilizes training. Our contributions can be summarized as follows:
@@
  • Extensive experiments on SD 1.5 sd15, SDXL sdxl, and the industrial-scale Ovis-U1 ovis-u1 show that our method consistently improves preference metrics while preserving or enhancing aesthetic quality, stabilizing training, and avoiding collapse. Moreover, the gains hold across text-to-image models, unified generators, and image-editing setups.
Figure 1: Training dynamics of preference losses during DPO finetuning without (left) and with (right) our safe-$\lambda$ mechanism on SD 1.5 sd15. Images beneath the plots illustrate samples generated at training steps $\{0,500,1000,1500,2000\}$.
2 Related Work.
@@
Diffusion models have become a leading paradigm for image synthesis, offering strong quality and diversity diffusion_survey. Denoising diffusion with a variational objective ddpm and continuous-time score-based formulations with SDEs gm_sde; diffusion_sde underpin modern systems. Refinements such as EDM edm and rectified flow or flow matching rectified_flow; flow_matching clarify objectives and improve robustness. Guidance-based conditioning cg; cfg enhances controllability. For text-to-image generation, latent diffusion sd15 enables efficient high-resolution synthesis and supports large systems like SD3 sd3 and FLUX flux. In parallel, unified generators handle text-to-image and image editing within a single model ovis-u1. Our method applies to both families and is architecture-agnostic, working with UNet unet-style and DiT dit-style backbones.
@@
Direct Preference Optimization llm_dpo; diffusion-dpo; nc-dpo has been adapted to diffusion models to align generation with human comparisons while avoiding full reinforcement learning. A broad class of variants lpo; mapo calibrates the preference margin or the relative branch influence to improve stability and protect the generation. Other approaches seek to guide the update directions and step magnitudes in LLMs llm_survey by employing subspace projections and modest objective clipping orthogonal-dpo; bounded-dpo; calibrated-dpo; robust-dpo; chi2-po. Related work such as DPOP dpop promotes positivity constraints to mitigate failure modes in preference optimization, and MaPPO mappo incorporates prior knowledge via a maximum-a-posteriori objective. Diffusion-specific methods further account for the multi-step nature of denoising by reweighting across timesteps or by adding entropy regularization, exemplified by Balanced-DPO balanceddpo, DSPO dspo, and SEE-DPO see-dpo. In contrast, our Diffusion-SDPO introduces a per-step, geometry-aware safe scaling factor based on the inner product between winner and loser output-space gradients, which provides direct control over the winner loss at each step while continuing to expand the preference margin.
@@ 3 Preliminaries
Diffusion models diffusion; ddpm construct a Markov chain that gradually corrupts clean data with additive noise and then learn a parametric denoiser to invert this corruption. Let a variance schedule $\{\beta_{t}\}_{t=1}^{T}$ be given and define $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. The forward process can be defined as:
@@
$\mathcal{L}_{\text{diffusion}}\;=\;\mathbb{E}_{\,x_{0}\sim p_{\text{data}},\;\,t\sim\mathrm{Uniform}\{1,\dots,T\},\;\,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\;\bigl\|\,\epsilon_{\theta}\!\bigl(x_{t},t\bigr)-\epsilon\bigr\|_{2}^{2}.$ (4)
Minimizing $\mathcal{L}_{\text{diffusion}}$ yields a time-aware denoiser that can be applied in reverse order to iteratively remove noise and synthesize new samples from an initial Gaussian latent.
@@
Given a prompt $c$ and two images $x^{w}_{0}$ (preferred, “winner”) and $x^{l}_{0}$ (less preferred, “loser”), preference alignment for diffusion models seeks parameters $\theta$ such that the model assigns higher likelihood to $x^{w}_{0}$ than to $x^{l}_{0}$ nc-dpo; diffusion-dpo. A diffusion sampler produces a trajectory $(x_{T},\ldots,x_{0})$ and, at each time $t$, a reverse conditional $p_{\theta}(x_{t}\mid x_{t+1},c)$ ddpm; diffusion; rectified_flow. To instantiate DPO in this setting, we adopt the standard formulation wherein the stepwise preference score is the log-likelihood ratio with respect to a frozen reference model diffusion-dpo; dspo:
@@
The Diffusion–DPO diffusion-dpo loss applies Bradley-Terry-style BT_model logistic regression to the winner-loser pair at the same $t$:
@@
$\mathcal{L}_{\text{Diffusion-DPO}}=-\;\mathbb{E}\!\left[\log\sigma\!\left(\beta\log\frac{p_{\theta}(x_{t}^{w}\!\mid x_{t+1}^{w},c)}{p_{\text{ref}}(x_{t}^{w}\!\mid x_{t+1}^{w},c)}-\beta\log\frac{p_{\theta}(x_{t}^{l}\!\mid x_{t+1}^{l},c)}{p_{\text{ref}}(x_{t}^{l}\!\mid x_{t+1}^{l},c)}\right)\right].$ (7)
Under common parameterizations, Eq. 5 reduces to simple residual comparisons. For DDPM-style Gaussians ddpm; diffusion, writing $\hat{\epsilon}_{\theta}=\epsilon_{\theta}(x_{t+1},c,t)$ for the predicted noise, $\hat{\epsilon}_{\text{ref}}=\epsilon_{\text{ref}}(x_{t+1},c,t)$ for the reference noise and $\epsilon$ for the ground-truth noise that forms $x_{t}$, the log-ratio can be expressed as:
@@
and an analogous expression holds for velocity or flow-matching parameterizations by replacing the noise residual with the corresponding target rectified_flow. For notational brevity, we write the stepwise contrastive objective as $\mathcal{L}(x_{t+1},c,t)=\frac{1}{2}\|\epsilon_{\theta}(x_{t+1},c,t)-\epsilon\|_{2}^{2}-\frac{1}{2}\|\epsilon_{\text{ref}}(x_{t+1},c,t)-\epsilon\|_{2}^{2}$. Hence, the winner and loser margin loss are defined as $\mathcal{L}^{w}=\mathcal{L}(x_{t+1}^{w},c,t)$ and $\mathcal{L}^{l}=\mathcal{L}(x_{t+1}^{l},c,t)$, respectively. Substituting Eq. 8 into Eq. 7 gives the training loss
@@
Substituting Eq. 8 into Eq. 7 yields an implementable objective whose inner term is the per-step error difference between winner and loser branches. Diffusion–DPO diffusion-dpo thus encourages decreasing the winner’s prediction error while increasing the loser’s at the same timestep. However, this objective does not guarantee a monotonic decrease of the winner loss. Empirically, over-penalizing the loser can also worsen the preferred sample. In the left part of Fig. 1, the margin $\mathcal{L}^{w}-\mathcal{L}^{l}$ becomes increasingly negative, yet both $\mathcal{L}^{w}$ and $\mathcal{L}^{l}$ increase, indicating degradation of absolute performance and potential instability or collapse. This exposes a gap between relative alignment (widening the margin) and absolute error control (preserving the preferred sample). The difficulty is that the winner and loser gradients are misaligned and vary across timesteps. We therefore introduce a simple stepwise update that, to first order, guarantees the preferred loss does not increase at each step while still promoting margin expansion.
@@ 4.1 Safe Update via First-Order Approximation
4.2 Closed-Form Safeguard in Output Space
Directly evaluating the parameter-space bound in Eq. 14 is infeasible for a high-dimensional model, since it would require computing and storing the full gradients $\nabla_{\theta}\mathcal{L}^{w}$ and $\nabla_{\theta}\mathcal{L}^{l}$ just to take their dot product. However, we can derive a convenient proxy by considering gradients in the model’s output space. Modern diffusion models predict a noise or image tensor as output, and the training loss (e.g., a denoising score-matching loss ddpm) is defined on this output. Let $o^{w}$ and $o^{l}$ denote the model’s output activations for the winner and loser branches respectively (for example, $o$ could be the predicted noise residual at a certain diffusion step). Using the chain rule, we have $\nabla_{\theta}\mathcal{L}^{w}={J^{w}}^{\top}\nabla_{o}\mathcal{L}^{w}$ and $\nabla_{\theta}\mathcal{L}^{l}={J^{l}}^{\top}\nabla_{o}\mathcal{L}^{l}$, where $J$ is the Jacobian $\partial o/\partial\theta$ and $\nabla_{o}\mathcal{L}$ is the gradient of the loss with respect to the model output. The term $\nabla_{\theta}{\mathcal{L}^{w}}^{\top}\nabla_{\theta}\mathcal{L}^{l}$ can then be written as:
$\nabla_{\theta}{\mathcal{L}^{w}}^{\top}\nabla_{\theta}\mathcal{L}^{l}=(\nabla_{o}\mathcal{L}^{w})^{\top}\big(J^{w}{J^{l}}^{\top}\big)(\nabla_{o}\mathcal{L}^{l}).$ (15)
If we assume a local near-isometry, namely $J^{w}{J^{l}}^{\top}\approx I$ (see Sec. A for more details), then the parameter-space inner product is well approximated by the Euclidean dot product in output space, $(\nabla_{o}\mathcal{L}^{w})^{\top}(\nabla_{o}\mathcal{L}^{l})$. Let $g^{w}=\nabla_{o}\mathcal{L}^{w}$ and $g^{l}=\nabla_{o}\mathcal{L}^{l}$ denote the output-space gradients for the winner and the loser. We then define the safe step-size as:
$\lambda_{\text{safe}}:=\frac{\|g^{w}\|_{2}^{2}}{{g^{w}}^{\top}g^{l}},$ (16)
if ${g^{w}}^{\top}g^{l}>0$, and we set $\lambda_{\text{safe}}=+\infty$ (i.e., impose no cap on $\lambda$) if ${g^{w}}^{\top}g^{l}\leq 0$. In words, $\lambda_{\text{safe}}$ is a closed-form upper bound on the loser weight based on output-space gradients. It is computed cheaply per batch by taking the dot product of the two model-output error signals. Whenever the loser’s error vector has a positive correlation with the winner’s error vector (${g^{w}}^{\top}g^{l}>0$), $\lambda_{\text{safe}}$ provides a finite limit to how strongly we can apply the loser’s gradient without risking an increase in the winner’s loss. On the other hand, if the errors are orthogonal or negatively correlated (${g^{w}}^{\top}g^{l}\leq 0$), the winner is not threatened by the loser’s update.
Input: DPO dataset $\mathcal{D}=\{(c,x_{0}^{w},x_{0}^{l})\}$; model $\epsilon_{\theta}$; reference $\epsilon_{\text{ref}}$; safety slack $\mu\!\in\![0,1]$; schedule length (timesteps) $T$; learning rate $\eta$.
@@ 4.2 Closed-Form Safeguard in Output Space
  2. Get $(x_{t+1}^{w},x_{t+1}^{l})$ from Eq. 2 and compute $\hat{\epsilon}^{w}_{\theta}=\epsilon_{\theta}(x_{t+1}^{w},c,t),\hat{\epsilon}^{l}_{\theta}=\epsilon_{\theta}(x_{t+1}^{l},c,t),\hat{\epsilon}^{w}_{\text{ref}}=\epsilon_{\text{ref}}(x_{t+1}^{w},c,t),\hat{\epsilon}^{l}_{\text{ref}}=\epsilon_{\text{ref}}(x_{t+1}^{l},c,t)$.

### v2

@@ 1 Introduction
Figure 1: Training dynamics of preference losses during DPO finetuning without (left) and with (right) our safe-$\lambda$ mechanism on SD 1.5 [sd15]. Images beneath the plots illustrate samples generated at training steps $\{0,500,1000,1500,2000\}$.
@@
Text-to-image diffusion models [diffusion_survey] have achieved remarkable success in generating diverse and high-quality images [flux, google2025_nano_banana]. However, aligning these powerful generative models with nuanced human preferences remains a critical challenge. Recent approaches have begun to incorporate human feedback [rlhf] into diffusion model training, drawing inspiration from alignment techniques used in large language models. In particular, Direct Preference Optimization (DPO) [llm_dpo] has emerged as a promising alternative to reinforcement learning for finetuning on human preferences. DPO directly optimizes the model on pairwise human comparisons (winner vs. loser outputs), and has been successfully adapted to text-to-image diffusion models in methods [diffusion-dpo, mapo, dspo, dmpo, diffusion-kto] to improve visual appeal and prompt alignment. Despite these advances, we find that existing DPO-based alignment of diffusion models still faces a fundamental limitation: simply maximizing the preference margin between “winner” and “loser” outputs does not necessarily translate to better absolute generation quality of the finetuned model.
@@
In our empirical analysis, we find that standard Diffusion-DPO [diffusion-dpo] exhibits unstable training dynamics, and the model’s generative quality can deteriorate as training proceeds. As illustrated in the left part of Fig. 1, we find that both the winner’s and loser’s denoising losses tend to increase over time, even though the preference margin ($\mathcal{L}^{w}-\mathcal{L}^{l}$) becomes more negative in the intended direction. This indicates that the model is widening the relative preference gap by making the less-preferred outputs worse, rather than truly improving the preferred outputs. In other words, relative alignment comes at the expense of absolute quality. The lack of a safeguard on the winner’s loss in existing DPO objectives leads to unstable training and potential collapse, corroborating observations in prior work [dpop, calibrated-dpo, see-dpo] that overly aggressive preference optimization can harm generative performance. These findings motivate the need for a new approach to preference-based diffusion finetuning that can increase preference alignment while preserving or improving the quality of the preferred outputs.
@@
To address this challenge, we propose Diffusion-SDPO [^1] Throughout the text, “Diffusion-SDPO” is used as a conceptual umbrella for our method and its guiding principles. When referring to concrete instantiations, we write “X+SDPO” to denote the integration of SDPO with a specific base DPO variant $X$ (e.g., Diffusion-DPO, DSPO, DMPO), which clarifies the application setting and configuration. – a Safeguarded Direct Preference Optimization method for diffusion models. The key idea in Diffusion-SDPO is to introduce a simple yet effective winner-preserving update rule that controls the influence of the loser sample’s gradient at each training step. In contrast to standard DPO [llm_dpo, diffusion-dpo] which updates the model by contrasting winner and loser equally, we derive an adaptive scaling factor for the loser’s gradient based on the geometry of the winner and loser gradients. Intuitively, our method downweights the loser branch’s contribution whenever its gradient is misaligned with the winner’s gradient. Grounded in a first-order analysis, the safeguard computes a closed-form $\lambda_{\text{safe}}$ from the inner product of the winner and loser gradients, guaranteeing that each step does not worsen the preferred output’s reconstruction loss. In practice, Diffusion-SDPO seamlessly modifies the DPO objective with this adaptive loser scaling (see Fig. 1, right), which expands the preference margin while strictly controlling the absolute error of preferred outputs. Notably, our approach is model-agnostic and can be applied on top of various diffusion alignment frameworks [diffusion-dpo, mapo, dmpo, dspo], acting as a plug-in optimizer that stabilizes training. Our contributions can be summarized as follows:
@@
  • Extensive experiments on SD 1.5 [sd15], SDXL [sdxl] (both UNets [unet]), and the industrial-scale Ovis-U1 [ovis-u1] (DiT [dit]) show that our method is architecture-agnostic. It delivers consistent improvements in preference metrics while preserving or enhancing aesthetic quality, stabilizing training, and avoiding collapse. These benefits persist across text-to-image, image editing and unified generation setups.
2 Related Work
@@
Diffusion models have become a leading paradigm for image synthesis, offering strong quality and diversity [diffusion_survey]. Denoising diffusion with a variational objective [ddpm] and continuous-time score-based formulations with SDEs [gm_sde, diffusion_sde] underpin modern systems. Refinements such as EDM [edm] and rectified flow or flow matching [rectified_flow, flow_matching] clarify objectives and improve robustness. Guidance-based conditioning [cg, cfg] enhances controllability. For text-to-image generation, latent diffusion [sd15] enables efficient high-resolution synthesis and supports large systems like SD3 [sd3] and FLUX [flux]. In parallel, unified generators handle text-to-image and image editing within a single model [ovis-u1]. Our method applies to both families and is architecture-agnostic, working with UNet [unet]-style and DiT [dit]-style backbones.
@@
Direct Preference Optimization [llm_dpo, diffusion-dpo, nc-dpo] has been adapted to diffusion models to align generation with human comparisons while avoiding full reinforcement learning. A broad class of variants [lpo, mapo] calibrates the preference margin or the relative branch influence to improve stability and protect the generation. Other approaches seek to guide the update directions and step magnitudes in LLMs [llm_survey] by employing subspace projections and modest objective clipping [orthogonal-dpo, bounded-dpo, calibrated-dpo, robust-dpo, chi2-po]. Related work such as DPOP [dpop] promotes positivity constraints to mitigate failure modes in preference optimization, and MaPPO [mappo] incorporates prior knowledge via a maximum-a-posteriori objective. Diffusion-specific methods further account for the multi-step nature of denoising by reweighting across timesteps or by adding entropy regularization, exemplified by Balanced-DPO [balanceddpo], DSPO [dspo], and SEE-DPO [see-dpo]. In contrast, our Diffusion-SDPO introduces a per-step, geometry-aware safe scaling factor based on the inner product between winner and loser output-space gradients, which provides direct control over the winner loss at each step while continuing to expand the preference margin.
@@ 3 Preliminaries
Diffusion models [diffusion, ddpm] construct a Markov chain that gradually corrupts clean data with additive noise and then learn a parametric denoiser to invert this corruption. Let a variance schedule $\{\beta_{t}\}_{t=1}^{T}$ be given and define $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. The forward process can be defined as:
@@
$\mathcal{L}_{\text{diffusion}}\;=\;\mathbb{E}_{x_{0},t,\epsilon}\;\bigl\|\,\epsilon_{\theta}\!\bigl(x_{t},t\bigr)-\epsilon\bigr\|_{2}^{2}.$ (4)
where $x_{0}\!\sim\!p_{\text{data}},t\!\sim\!\mathrm{Uniform}\{1,\dots,T\}$ and $\epsilon\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I})$. Minimizing $\mathcal{L}_{\text{diffusion}}$ yields a time-aware denoiser that can be applied in reverse order to iteratively remove noise and synthesize new samples from an initial Gaussian latent.
@@
Given a prompt $c$ and two images $x^{w}_{0}$ (preferred, “winner”) and $x^{l}_{0}$ (less preferred, “loser”), preference alignment for diffusion models seeks parameters $\theta$ such that the model assigns higher likelihood to $x^{w}_{0}$ than to $x^{l}_{0}$ [nc-dpo, diffusion-dpo]. A diffusion sampler produces a trajectory $(x_{T},\ldots,x_{0})$ and, at each time $t$, a reverse conditional $p_{\theta}(x_{t}\mid x_{t+1},c)$ [ddpm, diffusion, rectified_flow]. To instantiate DPO in this setting, we adopt the standard formulation wherein the stepwise preference score is the log-likelihood ratio with respect to a frozen reference model [diffusion-dpo, dspo]:
@@
The Diffusion–DPO [diffusion-dpo] loss applies Bradley-Terry-style [BT_model] logistic regression to the winner-loser pair at the same $t$:
@@
$\begin{split}\mathcal{L}_{\text{Diffusion-DPO}}=&-\,\mathbb{E}\!\Bigl[\log\sigma\Bigl(\beta\log\frac{p_{\theta}(x_{t}^{w}\mid x_{t+1}^{w},c)}{p_{\text{ref}}(x_{t}^{w}\mid x_{t+1}^{w},c)}\\ \qquad&-\beta\log\frac{p_{\theta}(x_{t}^{l}\mid x_{t+1}^{l},c)}{p_{\text{ref}}(x_{t}^{l}\mid x_{t+1}^{l},c)}\Bigr)\Bigr].\end{split}$ (7)
Under common parameterizations, Eq. 5 reduces to simple residual comparisons. For DDPM-style Gaussians [ddpm, diffusion], writing $\hat{\epsilon}_{\theta}=\epsilon_{\theta}(x_{t+1},c,t)$ for the predicted noise, $\hat{\epsilon}_{\text{ref}}=\epsilon_{\text{ref}}(x_{t+1},c,t)$ for the reference noise and $\epsilon$ for the ground-truth noise that forms $x_{t}$, the log-ratio can be expressed as:
@@
and an analogous expression holds for velocity or flow-matching parameterizations by replacing the noise residual with the corresponding target [rectified_flow]. For notational brevity, we write the stepwise contrastive objective as $\mathcal{L}(x_{t+1},c,t)=\frac{1}{2}\|\epsilon_{\theta}(x_{t+1},c,t)-\epsilon\|_{2}^{2}-\frac{1}{2}\|\epsilon_{\text{ref}}(x_{t+1},c,t)-\epsilon\|_{2}^{2}$. Hence, the winner and loser margin loss are defined as $\mathcal{L}^{w}=\mathcal{L}(x_{t+1}^{w},c,t)$ and $\mathcal{L}^{l}=\mathcal{L}(x_{t+1}^{l},c,t)$, respectively. Substituting Eq. 8 into Eq. 7 gives the training loss
@@
Substituting Eq. 8 into Eq. 7 yields an implementable objective whose inner term is the per-step error difference between winner and loser branches. Diffusion–DPO [diffusion-dpo] thus encourages decreasing the winner’s prediction error while increasing the loser’s at the same timestep. However, this objective does not guarantee a monotonic decrease of the winner loss. Empirically, over-penalizing the loser can also worsen the preferred sample. In the left part of Fig. 1, the margin $\mathcal{L}^{w}-\mathcal{L}^{l}$ becomes increasingly negative, yet both $\mathcal{L}^{w}$ and $\mathcal{L}^{l}$ increase, indicating degradation of absolute performance and potential instability or collapse. This exposes a gap between relative alignment (widening the margin) and absolute error control (preserving the preferred sample). The difficulty is that the winner and loser gradients are misaligned and vary across timesteps. We therefore introduce a simple stepwise update that, to first order, guarantees the preferred loss does not increase at each step while still promoting margin expansion.
@@ 4.1 Safe Update via First-Order Approximation
Input: Dataset $\mathcal{D}=\{(c,x_{0}^{w},x_{0}^{l})\}$; model $\epsilon_{\theta}$; reference $\epsilon_{\text{ref}}$; safety slack $\mu\!\in\![0,1]$; schedule length $T$; learning rate $\eta$.
@@
  2. Get $(x_{t+1}^{w},x_{t+1}^{l})$ from Eq. 2 and compute
  -6pt
$\hat{\epsilon}^{w}_{\theta}$ $=\epsilon_{\theta}(x_{t+1}^{w},c,t),$ $\hat{\epsilon}^{l}_{\theta}\,\;=\epsilon_{\theta}(x_{t+1}^{l},c,t),$
$\hat{\epsilon}^{w}_{\text{ref}}$ $=\epsilon_{\text{ref}}(x_{t+1}^{w},c,t),$ $\hat{\epsilon}^{l}_{\text{ref}}=\epsilon_{\text{ref}}(x_{t+1}^{l},c,t).$
  -18pt